# Fighting Exponentially Small Gaps by Counterdiabatic Driving

### v2

@@ III.1 Brief review of counterdiabatic driving
The adiabatic theorem of quantum mechanics guarantees that the quantum dynamics, generated by a time-dependent Hamiltonian $H[\lambda(t)]$, tracks its instantaneous eigenstates, provided that variations in time are slow enough. The adiabaticity condition reads
$\dot{\lambda}\,\frac{|\bra{m[\lambda(t)]}\partial_{\lambda}H[\lambda(t)]\ket{n[\lambda(t)]}|}{(E_{n}[\lambda(t)]-E_{m}[\lambda(t)])^{2}}\ll 1,$ (20)
where |$\ket{n[\lambda]}$ are the instantaneous eigenstates and $E_{n}[\lambda]$ the corresponding energies. If the condition above is met, then transitions between eigenstates are suppressed [52]; see as well Refs. [7, 8] for refinements on the adiabaticity condition. Yet, a perfect adiabatic tracking can be obtained at any finite driving rate, if the dynamics is generated by the modified counterdiabatic driving (CD) Hamiltonian [48, *Demirplak2005Assisted, *Demirplak2008Consistency, 51]
@@
where the last equation assumes that the spectrum of $H[\lambda(t)]$ is nondegenerate [51]. The first term on the right-hand side of $H_{1}[\lambda]$ is the generator of parallel transport familiar from proofs of the adiabatic theorem. The second term, being diagonal in the instantaneous eigenbasis, provides the correct Berry phase associated with a perfectly adiabatic trajectory [101].
The challenge in utilizing the CD Hamiltonian, such as the one in Eq. (22), comes from the fact that the terms involved are highly non-local and multiple-body, being |$\outerproduct{\partial_{\lambda}n}{n}$ | similar in structure to a projector [53]. For this reason, a growing body of works has been investigating the possibility of truncating the CD Hamiltonian to few-body local terms while retaining the suppression of diabatic transitions [53, 54, 56, 55, 58, 59, 76, 77]. A convenient framework to this end relies on the Floquet-Krylov expansion of the CD term [59, 60, 61]. Using the equivalent representation
[…]
@@
Here, $\alpha_{n}=(-1)^{n}\epsilon^{-2n}$ is formally diverging, but allowing the $\alpha_{n}$’s to be free variational parameters, one obtains a controlled expansion of the CD term with increasingly non-local operators [59]. To fix the parameters $\alpha_{n}$, one may either employ a QAOA-inspired procedure, in which they are updated recursively by using as cost function the energy at the end of the driving process [102, 62, 63, 64], or an a priori method like the minimization of the action [58]
@@
with respect to the parameters $\alpha_{n}$. We choose to follow the second route since it allows for greater analytical control and does not rely on heavy numerical optimization.
@@ III.2 Approximate counterdiabatic driving for the spin-glass bottleneck model
Following Refs. [58, 59], consider the lowest-order approximant to the CD term in the Floquet-Krylov expansion. At the first order, the commutator $[H,\partial_{\lambda}H]$, with $H$ in Eq. (1), contains terms $\sigma_{j}^{y}\sigma_{j+1}^{z}+\sigma_{j}^{z}\sigma_{j+1}^{y}$, motivating the variational ansatz
@@
The coefficients $\alpha_{j}$ can be found by minimizing the action (26). This yields the equations
$\left[8(1-\lambda)^{2}+\lambda^{2}\left(J_{j-1}^{2}+2J_{j}^{2}+J_{j+1}^{2}\right)\right]\alpha_{j}\\ +2\lambda^{2}J_{j}\left(J_{j-1}\alpha_{j-1}+J_{j+1}\alpha_{j+1}\right)=-J_{j},$ (29)
with periodic boundary conditions implied, which can be solved via linear algebra techniques with a cost polynomial in $L$. Going to the second order, from the computation of the nested commutator $[H,[H,[H,\partial_{\lambda}H]]]$, one determines the variational ansatz $H_{1}^{(1)}[\lambda]+H_{1}^{(2)}[\lambda]$, where
@@
Again, the coefficients $\{\alpha_{j},\beta_{j}\}$ can be determined by solving a $2L\times 2L$ linear system (Eq. (125)). Generalizing, one needs to solve a $nN\times nN$ linear system for the $n$-th order term of the Floquet-Krylov expansion. Thus, for finite $n$, the CD term can be determined efficiently with the help of a classical computer.
@@ IV.1 Counterdiabatic driving and gap amplification
According to the adiabatic theorem, adiabaticity can be preserved upon reducing the driving time, provided that the gap is increased by the same factor. However, the construction of the CD term outlined in Sec. III.1 indicates that CD is more than a mere gap amplification method: the operators involved are constructed in such a way to transport a state along its adiabatic path, and they do so by enforcing parallel transport while accounting the corresponding quantum phase. By contrast, catalysts and gap-amplifying terms  just enhance the dephasing due to highly oscillatory terms $e^{i(E_{m}-E_{n})t}$ for $m\neq n$, making the adiabatic approximation hold to a better degree [30, 31, 32, 33, 34, 35, 36, 37, 38, 39].
@@ IV.2 Performance of counterdiabatic driving
Above, the adiabatic theorem was used to predict the performance of approximate CD. It was found that the gap remains exponentially small in the system size for all driving times, with the gain of CD being less and less significant as the driving time increases. In order to quantify precisely to what extent CD helps the driving process, however, we next integrate the time-dependent Schrödinger equation and see whether the success probability is actually increased.
As before, a cubic ramp $\lambda(t)=3(t/T)^{2}-2(t/T)^{3}$, satisfying $\dot{\lambda}(0)=\dot{\lambda}(T)=0$, is employed. Upon initializing the system in the paramagnetic ground state of $H[\lambda=0]$, we consider the evolution with the approximate CD Hamiltonians $H[\lambda(t)]+\dot{\lambda}(t)H^{(1)}[\lambda(t)]$ (first order ansatz) or $H[\lambda(t)]+\dot{\lambda}(t)[H^{(1)}[\lambda(t)]+H^{(2)}[\lambda(t)]]$ (second order ansatz). Because of the free-fermionic nature of the model under consideration, we evolve the single-particle operators $(c_{j}^{\phantom{\dagger}},c_{j}^{\dagger})$ or $\Gamma_{k}^{\pm}$ by means of Bogoliubov-de Gennes equations, and then compute the desired expectation values at the end of the protocol with the appropriate contractions; see App. E for all the details. This technique allows one to reach much larger system sizes, which are needed in the present case to extract the asymptotic scalings. Incidentally, this is the reason why we employ the chain in Eq. (1) instead of examples of NP-hard Ising Hamiltonians [103].
@@
The results of our numerical simulations to assess the performance of the driving are presented in Figs. 7. First, we compute the kink number in Fig. 7, i.e., the number of frustrated bonds at the end of the process. It is given by the expectation value on the final state of the operator
@@
The findings of Fig. 7 agree with the known results about the Kibble-Zurek (KZ) density of defects formed upon crossing a quantum phase transition [79, 80, 81, 82, 83]. The collapse in Fig. 7b shows that the defect density scales as $\sim T^{-\nu/(1+z\nu)}=T^{-1/2}$, where $\nu=1$ is the correlation length exponent of the transverse-field Ising model and $z=1$ is its dynamical exponent. The presence of the three weak links manifests only in the breakdown of the KZ scaling at large times, with the formation of the plateau. The breakdown of the KZ scaling at short times, due to the introduction of approximate CD terms, confirms the early results of Ref. [53] in this setting. Thus, our findings demonstrate that CD improves fast driving at short times but does not help significantly to overcome the exponentially small gap, as further discussed in Sec. VI.
@@ V Quantum Brachistochrone Counterdiabatic Driving (QBCD): exponential speedup
In this section, we show how the efficiency of the driving can be improved beyond the low-order CD expansions by canceling the transitions between the ground state and the first excited state around the critical point $\lambda=\lambda_{c}$.
@@ V.1 Comparison to the local approach in the spin glass toy model
Above, the first- and second-order ansatzes were used without requiring any knowledge of the instantaneous spectral properties. Here, an approximate knowledge of the spectrum at a single point in parameter space is assumed, and much better results are obtained. This approach is reminiscent of reverse quantum annealing, where an approximate ground state of the problem Hamiltonian is used to improve the performance of the adiabatic algorithm [104, 105, 106]. In addition, this method differs from the early approaches to CD in many-body systems, where the exact CD term was truncated according to the locality of the operators involved [53, 56]. Here, instead, the complete CD Hamiltonian is employed, but with an approximate knowledge of the eigenvalues and eigenvectors. The truncation is, in turn, performed in the number of parameter values at which it is evaluated. Although in the worst-case scenario, this does not reduce the computational cost of the search for the final ground state when compared to diagonalizing the final Hamiltonian, the use of additional numerical methods such as reverse annealing or quantum Monte Carlo to extract approximate knowledge of the critical point are feasible strategies to keep the numerical cost under control. The simplest version of this approach is to take a single parameter value at which the approximate knowledge is required.
The idea is to focus on the bottleneck of the driving dynamics and use approximate left- and right-localized edge states, which are responsible for the exponentially closing gap, to construct the CD term at the critical point. By multiplying such a term with the time-dependent derivative of the schedule, it is ensured that the spectrum is not altered at the endpoints $\lambda=0$ and 1 (see App. C.4). In the language of the one-particle effective model, Eqs. (16)–(17), the QBCD Hamiltonian reads
@@
where $\lambda_{c}$ denotes the point in parameter space where the gap is minimal and equal to $\Delta_{\min}$, $|\psi_{\mathrm{L,R}})$ denote the left/right localized edge states in the $L$-dimensional linear space of the effective chain. Equation (35), inspired by CD, shares the structure of the Hamiltonian in the quantum brachistochrone problem for the time-optimal evolution within this two-level subspace [107, 108, 109], motivating the term QBCD. Upon contracting with the $\Gamma$ operators, one obtains the full expression of the QBCD Hamiltonian. The gap and the left and right eigenvectors are obtained via analytical approximations valid in the close neighborhood of the critical point (see App. C.4 Notice also that $\mathcal{H}^{\mathrm{QBCD}}$ takes the form of a $\sigma^{y}$ Pauli matrix in the reduced subspace of the left- and right-localized single-particle states.
Figure 8: Gap amplification from the QBCD term, Eq. (35). Comparing with Fig. 5, one can see that $H^{\mathrm{QBCD}}$ is able to enlarge the minimal gap on the adiabatic path by a significant amount for all system sizes and driving times.
The QBCD Hamiltonian, Eq. (35), cancels exactly the transition between the ground and first excited states only at the gap-closing point; however, it also provides an efficient CD term in the neighborhood thereof. This feature can be inferred from the results shown in Fig. 8 and 9: both the gap $\Delta_{{\min},\mathrm{QBCD}}$ and the ground state fidelity are significantly increased with respect to the bare Hamiltonian. Here, the fidelity is captured by the kink number, since a value $K<1$ can come only from the ground state. In particular, one can see from Fig. 9a that for values of $T$ for which the bare Hamiltonian exhibits a kink number plateauing very close to 1, the QBCD term is able to lower it to $\expectationvalue{K}_{T}\approx 0.5$. This implies that the ground state is found at the end of the protocol with a finite probability of $\approx 0.5$, which is significantly larger than the one found with the bare protocol.
The effects of QBCD remain remarkably robust against deviations from the true critical point $\lambda_{c}$. This robustness is especially valuable in real-world applications, where experimental noise and fluctuations often hinder the precise identification of critical parameters. To test this, we studied the behavior of the minimal gap by varying the estimated critical point $\lambda^{*}_{c}$ within the range $\lambda_{c}\in[0.85\lambda^{*}_{c},1.15\lambda^{*}_{c}]$.
Beyond analytical approximations (App. C.4), we extended our analysis by performing a fully numerical search for the ground state energy gap and corresponding eigenstates using only 40 sampling points. This approach requires significantly less computational effort than simulating the full QBCD protocol, yet still captures the essential features needed for its construction. As such, it broadens its applicability to a wide class of spin-glass systems, even in the absence of analytical control over the critical properties. Figure 10 shows the performance of the resulting QBCD Hamiltonian, constructed from the corresponding eigenstates $\left\lvert\psi_{R}\right)$ and $\left\lvert\psi_{L}\right)$, and the numerically determined minimal gap $\Delta_{\mathrm{min,QBCD}}$ at the shifted $\lambda_{c}$. Apart from verifying that the QBCD can achieve a polynomial decreasing gap with respect to $L$, this scaling also remains stable up to a constant factor across deviations up to approximately $10\%$ from the true critical point.
Interestingly, for the analytically approximated QBCD, the minimal gap even increases slightly with small shifts away from the critical point. Although the gap eventually decreases—for example, at $\lambda_{c}=1.125,\lambda_{c}^{*}$—the numerically obtained QBCD Hamiltonian still achieves a two-order-of-magnitude improvement over local expansions for system sizes as large as $L=600$. Constructing the QBCD using approximate numerical or experimental methods that demand far less computational effort than running the full simulation, thus suffices to realize exponential speedups in adiabatic protocols.
@@
Figure 10: Robustness of the minimal gap scaling against deviations between the applied QBCD Hamiltonian and that taken exactly at $\lambda_{c}$. Red dashed lines correspond to the approximate analytical results, while the blue dotted ones follow from a pure numerical approach. The black solid line corresponds to the analytical approach at $\lambda_{c}=\lambda^{*}_{c}$. The curves clearly indicate a polynomial decay rather than an exponential one.
Finally, let us investigate the efficiency of the variational and QBCD approaches presented from the perspective of energy resources. While reducing external resources is essential for experimental realizations [110, 111], the energetic costs determine the speed at which the assisted process can stay adiabatic [50, 112, 113]. The Hilbert-Schmidt norm of the Hamiltonians, defined as $\|H_{1}\|^{2}=\mathrm{Tr}[H^{\dagger}_{1}\,H_{1}]$, provides a natural measure to quantify the cost of CD [50, 53, 60, 61]. It can be shown (see App. D) that the energy costs of the low-order expansions grow as
@@
for the first- and second-order variational expansions, respectively. The QB approach surpasses the variational method in this aspect as well. As shown in App. D, the numerator and the denominator in Eq. (35) grow as (to leading order in $\ell$)
$\Delta_{\min}\simeq\ell(\lambda-1)e^{-(\ell-1)\kappa}(1+e^{-2\mu})\\ +\ell\lambda\left(e^{-\ell\kappa-2\mu}+e^{-(\ell-2)\kappa}\right)$ (37)
and
$(\psi_{\mathrm{R}}|\partial_{\lambda}\mathcal{H}[\lambda_{c}]|\psi_{\mathrm{L}})\approx\\ \ell\left(e^{-(\ell-1)\kappa}(1+e^{-2\mu})+e^{-\ell\kappa-2\mu}+e^{-(\ell-2)\kappa}\right).$ (38)
The two equations above imply a constant norm to leading order:
@@
To also incorporate the time dependence in the computation, in Fig. 11 the quantities $\frac{1}{T}\int_{0}^{T}\mathrm{d}t\,\|H_{1}\|^{2}$ are presented. It is clearly demonstrated that $\|H_{1}^{(1)}\|$ and $\|H_{1}^{(2)}\|$ increase with the system size, while $\|\mathcal{H}^{\mathrm{QBCD}}\|$ remains independent of $L$ to leading order. Thus, the QBCD approach not only achieves significantly larger fidelities to the target state, but it does so with a constant energy cost.
@@
Let us stress that the specific features of the model and its free-fermionic representation were used solely to access large system sizes, to compare the purely numerical construction with analytical approximations, and to also test analytically its energetic costs, while the general effects of the QBCD are expected to hold across a wide range of spin glass problems as detailed in the following sections.

### v3

@@ III.1 Brief review of counterdiabatic driving
The adiabatic theorem in quantum mechanics states that the time-evolved state of a system driven by a slowly varying Hamiltonian $H[\lambda(t)]$ closely follows the corresponding instantaneous eigenstate. A sufficient condition for adiabaticity is
[…]
@@
where the last equation assumes that the spectrum of $H[\lambda(t)]$ is nondegenerate [51]. The first term on the right-hand side of $H_{1}[\lambda]$ is the generator of parallel transport, familiar from proofs of the adiabatic theorem. The second term, which is diagonal in the instantaneous eigenbasis, provides the correct Berry phase associated with a perfectly adiabatic trajectory [100].
The challenge in utilizing the CD Hamiltonian, such as the one in Eq. (15), comes from the fact that the terms involved are highly non-local and multiple-body, being $|\outerproduct{\partial_{\lambda}n}{n}$ similar in structure to a projector [53]. For this reason, a growing body of works has investigated the possibility of truncating the CD Hamiltonian to few-body local terms, while retaining the suppression of diabatic transitions [53, 54, 56, 55, 58, 59, 76, 77]. A convenient framework for this purpose relies on the Floquet-Krylov expansion of the CD term [59, 60, 61]. Using the equivalent representation
$H_{1}[\lambda]=-\int_{0}^{+\infty}dx\,e^{-\epsilon|x|}\,e^{ixH[\lambda]}\partial_{\lambda}H[\lambda]e^{-ixH[\lambda]},$ (17)
@@
Here, $\alpha_{n}=(-1)^{n}\epsilon^{-2n}$ is formally diverging, but allowing the $\alpha_{n}$’s to be free variational parameters, one obtains a controlled expansion of the CD term with increasingly non-local operators [59]. To fix the parameters $\alpha_{n}$, one may employ a QAOA-inspired procedure, in which they are updated recursively by using as a cost function the energy at the end of the driving process [101, 62, 63, 64], or an a priori method like the minimization of the action [58]
@@
with respect to the parameters $\alpha_{n}$. We choose to follow the second route, since it allows greater analytical control and does not rely on heavy numerical optimization.
@@ III.2 Approximate counterdiabatic driving for the spin-glass bottleneck model
Following Refs. [58, 59], consider the lowest-order approximant to the CD term in the Floquet-Krylov expansion. At the first order, the commutator $H,\partial_{\lambda}H]$, with $H$ in Eq. (1), contains terms $\sigma_{j}^{y}\sigma_{j+1}^{z}+\sigma_{j}^{z}\sigma_{j+1}^{y}$, motivating the variational ansatz
@@
The coefficients $\alpha_{j}$ can be found by minimizing the action (19), as given by Eq. (112) in App. C.2. Going to the second order, from the computation of the nested commutator $H,[H,[H,\partial_{\lambda}H]]]$, one determines the variational ansatz $H_{1}^{(1)}[\lambda]+H_{1}^{(2)}[\lambda]$, where
@@
Again, the coefficients $\{\alpha_{j},\beta_{j}\}$ can be determined by solving a $2L\times 2L$ linear system of equations (Eq. (113)). Generalizing, one needs to solve a $nL\times nL$ linear system for the $n$-th order term of the Floquet-Krylov expansion. Thus, for finite $n$, the CD term can be determined efficiently with the help of a classical computer.
@@ IV.1 Counterdiabatic driving and gap amplification
According to the adiabatic theorem, adiabaticity can be preserved upon reducing the driving time, provided that the gap is increased by the same factor. However, the construction of the CD term outlined in Sec. III.1 indicates that CD is more than a mere gap amplification method: the operators involved are constructed in such a way to transport a state along its adiabatic path, and they do so by enforcing parallel transport while accounting for the corresponding quantum phase. By contrast, catalysts and gap-amplifying terms  just enhance the dephasing due to highly oscillatory terms $e^{i(E_{m}-E_{n})t}$ for $m\neq n$, making the adiabatic approximation hold to a better degree [30, 31, 32, 33, 34, 35, 36, 37, 38, 39].
@@ IV.2 Performance of counterdiabatic driving
Above, the adiabatic theorem was used to predict the performance of approximate CD. It was found that the gap remains exponentially small in the system size, with the relative advantage of CD decreasing for longer protocols. To quantify more precisely how much CD improves the adiabatic evolution, we next consider the average number of bonds violated compared to the final ground state.
As before, a cubic ramp $\lambda(t)=3(t/T)^{2}-2(t/T)^{3}$, satisfying $\dot{\lambda}(0)=\dot{\lambda}(T)=0$, is employed. Upon initializing the system in the paramagnetic ground state of $H[\lambda=0]$, we consider the evolution with the approximate CD Hamiltonians $H[\lambda(t)]+\dot{\lambda}(t)H^{(1)}[\lambda(t)]$ (first-order ansatz) or $H[\lambda(t)]+\dot{\lambda}(t)[H^{(1)}[\lambda(t)]+H^{(2)}[\lambda(t)]]$ (second-order ansatz). Because of the free-fermionic nature of the model under consideration, we evolve the single-particle operators $(c_{j}^{\phantom{\dagger}},c_{j}^{\dagger})$ or $\Gamma_{k}^{\pm}$ by means of Bogoliubov-de Gennes equations, and then compute the desired expectation values at the end of the protocol with the appropriate contractions; see App. E for all the details. This technique allows one to reach much larger system sizes, which are needed in the present case to extract the asymptotic scalings.
@@
The results of our numerical simulations to assess the performance of the driving are presented in Figs. 6. First, we compute the kink number in Fig. 6, i.e., the number of frustrated bonds at the end of the process. It is given by the expectation value of the operator on the final state
@@
The findings of Fig. 6 are consistent with the Kibble-Zurek (KZ) prediction for the density of defects formed in a quantum phase transition [79, 80, 81, 82, 83]. The collapse in Fig. 6b confirms the expected scaling $\sim T^{-\nu/(1+z\nu)}=T^{-1/2}$, with $\nu=1$ and $z=1$ for the transverse-field Ising model. The three weak links affect the dynamics only at long times, where the KZ scaling gives way to a plateau for the defect density. For fast driving protocols, deviations induced by approximate CD terms reproduce the results of Ref. [53]. Overall, CD improves the performance in the fast-driving regime but remains ineffective in mitigating excitations across the exponentially small gap, as discussed in Sec. VI.
@@ V Quantum Brachistochrone Counterdiabatic Driving (QBCD): exponential speedup
In this section, we show how the efficiency of the driving can be improved beyond low-order CD expansions by canceling ground–first-excited transitions near the critical point $\lambda=\lambda_{c}$.
@@ V.1 Comparison to the local approach in the spin glass toy model
While earlier ansatze required no spectral information, here we assume approximate knowledge of the spectrum at a single-parameter point. This resembles reverse quantum annealing, where an approximate ground state improves adiabatic performance [102, 103, 104]. Unlike many early many-body CD approaches based on truncating nonlocal terms [53, 56], which can be as costly as full diagonalization, methods like reverse annealing [102, 103, 104] or quantum Monte Carlo methods [1, 2, 105] can provide the needed approximate spectral data at moderate cost.
The key idea is to target the bottleneck by using approximate left and right localized edge states, responsible for the exponentially small gap, to construct the CD term at the critical point. Multiplying it by the schedule derivative preserves the spectrum at $\lambda=0,1$ (see App.C.4). In the one-particle effective model, Eq.(9), and Eq.(10), the QBCD Hamiltonian is
@@
[…]
Figure 7: Gap amplification from the QBCD term, Eq. (27). Comparing with Fig. 4, one can see that $H^{\mathrm{QBCD}}$ is able to significantly enlarge the minimal gap on the adiabatic path for all system sizes and driving times.
The QBCD Hamiltonian, Eq. (27), cancels exactly the transition between the ground and first excited states only at the gap-closing point; however, it also provides an efficient CD term in the neighborhood thereof. This feature can be inferred from the results shown in Figs. 7 and 8: both the gap $\Delta_{{\min},\mathrm{QBCD}}$ and the ground-state fidelity are significantly increased with respect to the bare Hamiltonian. Here, the fidelity is captured by the kink number, since a value $K<1$ can arise only from the ground state. In particular, from Fig. 8a one can see that for values of $T$ for which the bare Hamiltonian exhibits a kink number plateauing very close to 1, the QBCD term is able to lower it to $\expectationvalue{K}_{T}\approx 0.5$. This implies that the ground state is found at the end of the protocol with a finite probability of $\approx 0.5$, which is significantly greater than the one found with the bare protocol.
@@
The effects of QBCD remain remarkably robust against deviations from the exact critical point $\lambda_{c}$, an essential feature in realistic settings where noise and fluctuations hinder precise parameter identification and control. Furthermore, this robustness reduces computational costs in classical simulations and provides a resource-efficient route to approximate local CD protocols on quantum hardware, particularly when suitable truncations are applied [109, 110].
To demonstrate this, we vary the analytically obtained critical point $\lambda_{c}^{*}$ (App. C.4) within $0.85\lambda_{c}^{*},1.15\lambda_{c}^{*}]$ and study the behavior of the minimal gap. We also perform a purely numerical search for the ground state gap and eigenstates using only $40$ sampling points and vary $\lambda_{c}$ within the same interval in this approach. This requires far less computational effort than simulating the full QBCD protocol, yet captures the ingredients needed for its construction. As shown in Fig. 9, the gap at the critical point preserves its polynomial decay with the system size $L$ up to $\sim 10\%$ deviations from $\lambda_{c}$. Interestingly, for analytically approximated QBCD, the minimal gap can even increase slightly under small shifts, while for larger deviations (e.g., $\lambda_{c}=1.125\lambda_{c}^{*}$), the numerical construction still outperforms local expansions by two orders of magnitude for system sizes up to $L=600$. Thus, constructing QBCD from approximate spectral data is sufficient to realize an exponential speedup of adiabatic time scales across an exponentially small gap.
Figure 9: Robustness of the minimal gap scaling against deviations between the applied QBCD Hamiltonian and that taken exactly at $\lambda_{c}$. The red dashed lines correspond to the approximate analytical results, while the blue dotted lines follow from a purely numerical approach. The black solid line corresponds to the analytical approach at $\lambda_{c}=\lambda^{*}_{c}$. The curves clearly indicate a polynomial decay rather than an exponential one.
Finally, let us investigate the efficiency of the variational and QBCD approaches presented from the perspective of energy resources. Although reducing external resources is essential for experimental realizations [111, 112], the energetic costs determine the speed at which the assisted process can stay adiabatic [50, 113, 114]. The Hilbert-Schmidt norm of the Hamiltonians, defined as $\|H_{1}\|^{2}=\mathrm{Tr}[H^{\dagger}_{1}\,H_{1}]$, provides a natural measure to quantify the cost of CD [50, 53, 60, 61]. It can be shown (see App. D) that the energy costs of the low-order expansions grow as
@@
for the first- and second-order variational expansions, respectively. The QBCD exhibits more favorable features (App. D), as its norm is bounded in the leading order with $L$,
@@
To also incorporate the time dependence in the computation, in Fig. 10, the quantities $\frac{1}{T}\int_{0}^{T}\mathrm{d}t\,\|H_{1}\|^{2}$ are presented. It is clearly demonstrated that $\|H_{1}^{(1)}\|$ and $\|H_{1}^{(2)}\|$ increase with the system size, while $\|\mathcal{H}^{\mathrm{QBCD}}\|$ remains independent of $L$ up to the leading order. Thus, not only does the QBCD approach achieve significantly larger fidelities to the target state, but it does so in a resource-efficient way.
@@
We stress that the model’s specific features and free-fermion representation were used primarily to access large system sizes, compare numerical constructions with analytical approximations, and evaluate energetic costs. As we show in the following subsection, QBCD retains its efficiency in more realistic spin-glass models.